OpenStudy (anonymous):

Graph the two equations to solve the system. What is the solution? 3x + 2y = 6 and y = -3/2 x - 1 (1, 1.5) (2, 0) no solution infinitely many solutions

OpenStudy (zpupster):

here is the graph from desmos

OpenStudy (zpupster):

we can see there is no intersection to the lines --meaning no solutions. parallel lines have same slope y=mx+b m is the slope y= -3/2x - 1 -3/2 is the slope if we change the standard line form equation Ax+Bx = C into slope intercept form y=mx+b 3x + 2y = 6 3x-3x+2y = 6-3x 2y = 6-3x 2y/2 = 6-3x/2 y = -3/2 x +3 slope - 3/2 same slope = parallel therfore no solution.